OpenStudy (anonymous):

If f(x) = -5x + 1, find f(-3). Answer -14 -7 15 16 Just say a b c or d and I will award you a medal.

OpenStudy (anonymous):

d

OpenStudy (anonymous):

D! f(-3): -5*(-3) + 1 = 15 + 1 = 16